OpenStudy (toxicsugar22):

According to the following reaction, how many moles of water will be formed upon the complete reaction of 28.7 grams of hydrogen sulfide with excess oxygen gas? 2 H2S (g) + 3 O2 (g) = 2 H2O (l) + 2 SO2 (g)

OpenStudy (anonymous):

molesH2S=28.7/MMH2S, that is 844.11764*10^-3 2H2S=3O2,THEN 844.11764*10^-3=XO2,WHICH IS 1.2661764 MASS OF H2O=MOLES*MM=22.791176 Grams

OpenStudy (anonymous):

still no need to search for limiting reagent nut cracking so we ll just dive into our work the same drill again as we ve been doing let me check if the reaction is balanced, good 1H2O=1ZNOH2,THEN XMOLE OF H2O=0.446MOLES OF ZHOH2,WHICH IS 0.446MOLES OF H2O NOW moles=mass/MM,therefore mass of H2O=0.446*18=8.028 Grams

OpenStudy (anonymous):

absolutely

OpenStudy (anonymous):

dont worry you ll get it under your grip and it ll start begging you

OpenStudy (toxicsugar22):

According to the following reaction, how many moles of iron(III) chloride will be formed upon the complete reaction of 22.7 grams of iron with excess chlorine gas? 2 Fe (s) + 3 Cl2 (g) = 2 FeCl3 (s)

OpenStudy (anonymous):

just look at it again and ask me questions on the solution that might not seem too clear

OpenStudy (anonymous):

sorry give me a minute messed up a little here

OpenStudy (toxicsugar22):

ok

OpenStudy (anonymous):

405.3571428 * 10^-3 * (56+35.5*3)=65.870535

OpenStudy (toxicsugar22):

65.870535 that is the answer

OpenStudy (anonymous):

65.870535

OpenStudy (anonymous):

did you get it

OpenStudy (toxicsugar22):

kind of

OpenStudy (toxicsugar22):

I am getting it

OpenStudy (anonymous):

let us go through our drill this time the limiting reagent is Fe as the question says "with excess chlorine gas" ,also the question gives you a lead by giving you only the mass of Fe now moles of FE=mass/MM,which is 22.7/56 and value is 405.3571428 * 10^-3 from reaction if 2 Fe=2 FeCl,then 405.3571428 * 10^-3=X moles of FeCl which means moles of FeCl= 405.3571428 * 10^-3 Moles of FeCl=mass/MM therefore mass= 405.3571428 * 10^-3 * (56+35.5*3)= 65.870535 Grams

OpenStudy (toxicsugar22):

but it says inthe question "how many moles" of iron(III) chloride will be formed upon the complete reaction of 22.7 grams of iron with excess chlorine gas?

OpenStudy (toxicsugar22):

2 Fe (s) + 3 Cl2 (g) = 2 FeCl3 (s)

OpenStudy (anonymous):

sorry i think i am getting tired the simple answer is 405.3571428 * 10^-3 Moles

OpenStudy (anonymous):

let us go through our drill this time the limiting reagent is Fe as the question says "with excess chlorine gas" ,also the question gives you a lead by giving you only the mass of Fe now moles of FE=mass/MM,which is 22.7/56 and value is 405.3571428 * 10^-3 from reaction if 2 Fe=2 FeCl ,then 405.3571428 * 10^-3=X moles of FeCl which means moles of FeCl3= 405.3571428 * 10^-3

OpenStudy (toxicsugar22):

i got .4053571428

OpenStudy (anonymous):

excellent Great JOB way to go

OpenStudy (anonymous):

.4053571428=405.3571428 * 10^-3 ARE THE SAME THING JUST I USED MY ANDROID CALCULATOR

OpenStudy (toxicsugar22):

is this the anwere in moles

OpenStudy (anonymous):

YES IN MOLES

OpenStudy (toxicsugar22):

thank you